# Open-Source Reinforcement Learning Environments Implemented in MuJoCo with Franka Manipulator

### v2

@@ V CONCLUSION AND FUTURE WORK
Three open-source environments corresponding to three manipulation tasks, FrankaPush, FrankaSlide, and FrankaPickAndPlace, are proposed in this paper, where each task follows the Multi-Goal Reinforcement Learning framework. The environments are built with the Franka Emika Panda arm in MuJoCo Menagerie and feasible simulation attributes. Three representative off-policy algorithms with Hindsight Experience Replay are implemented to validate the feasibility of each environment. Benchmark results show that additional tasks that depend on MuJoCo and Franka arm can be developed by referring to the constructed environments in this paper.
Future work will further dive into tuning the MuJoCo simulation options to be as faithful as possible to the real environment. Extra tasks and manipulators with different end-effectors can also be considered to be added to the current work.

### v3

@@ V CONCLUSION AND FUTURE WORK
Three open-source environments corresponding to three manipulation tasks are proposed in this paper, where each environment follows the Multi-Goal Reinforcement Learning framework. These benchmark environments are built with the Franka model in MuJoCo Menagerie and powered by MuJoCo. Note that each environment is defined in a clean way in which unnecessary dependencies and modules are omitted. Three representative off-policy algorithms with Hindsight Experience Replay are implemented to validate the feasibility of each environment. To the best of our knowledge, there are a few benchmark environments built on top of MuJoCo Menagerie. We expect that our open-source environments can gain some attention to create more benchmark environments with MuJoCo Menagerie and thus lower barriers for researchers to test algorithms.
However, there are some limitations in our work. More diverse and realistic controllers, like the impedance controller, should be adopted to minimize the sim2real gap. Moreover, a smaller timestep has not been tested in MuJoCo considering the computing resources that we have. Future work will further dive into these limitations to be as faithful as possible to the real environment and offer more test results. Extra tasks and manipulators can also be considered to be added to the current work.